The product of two positive consecutive integers is 41 more than their sum. Find the integers.
Answer by checkley77(12844)

Let x & x+1 be the 2 integers.
x(x+1)=x+x+1+41
x^2+x=2x+42
x^2+x-2x-42=0
x^2-x-42=0
(x-7)(x+6)=0
x-7=0
x=7 ans. for the smaller integer.
7=1=8 ans. for the larger integer.
Proof:
7*8=7+8+41
56=56